Newton’s law of cooling assumes that air at room temperature is blown past the cooling body (forced cooling). For cooling in still air (natural cooling) a better modal is to assume that the rate of temperature decrease of the cooling body is directly proportional to the the (5/4)th power of the difference between the temperature u of the body and the temperature s of the surrounding air.
i) Write the law for natural cooling as a differential equation. Is this equation linear?
ii) Solve the equation obtained in i) above assuming that initially, the temperature of the cooling
body was u0.
1
Expert's answer
2020-03-10T11:19:49-0400

"-du\/dt=k(u-s)^{5\/4}"

is the differential equation for law of natural cooling.

Now, as degree of the differential equation is 1, thus it is linear.

"\int_{u_0}^u du/(u-s)^{5/4}=-\int_0^t k dt"

"(u-s)^{-1\/4} -(u_0 -s)^{-1\/4}=-4kt"

"u=s+((u_0 -s)^{-1\/4}-4kt)^{-4}"

is the solution for the above differential equation.